OpenStudy (anonymous):

cos2theta + 2cos^2theta=2

OpenStudy (anonymous):

kno any thing about square root

myininaya (myininaya):

so we have 3cos^2theta=2

myininaya (myininaya):

cos^2thetha=2/3

myininaya (myininaya):

take the square root of both sides costheta=2/3 and costhetha=-2/3

myininaya (myininaya):

thetha=arccos(2/3) and thetha=arcsos(-2/3)

OpenStudy (anonymous):

how did you get 3cos^2theta=2 if they aren't like terms so you can't combine them

myininaya (myininaya):

oh I thought they were both raised to the second power my bad

myininaya (myininaya):

i misread it

OpenStudy (anonymous):

oh ok but do you understand it?

myininaya (myininaya):

yes

myininaya (myininaya):

give me just a sec

myininaya (myininaya):

scanning right now

OpenStudy (anonymous):

thanks

myininaya (myininaya):

so it looks good?

OpenStudy (anonymous):

yea, i got the same answer before doing it a different longer complex way but its good.

myininaya (myininaya):

k